# Diffusion Models for Computational Design at the Example of Floor Plans

### v1

@@ 2.5 Fine-tuning Methodology
In order to fine-tune Stable Diffusion to the different semantic encodings, we implemented an algorithmic design approach in Python that generates a desired number of floor plans and then encodes it in each of the four different proposed styles $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ as well as a description prompt for each image. This ensures that all four models are on even ground during the training phase. Fig. 6 shows example images for each of the proposed encodings.
@@
The generated prompts describe the generated floor plan and follows the common guidelines for Stable Diffusion prompts, specifically for automatic1111: Individual concepts are separated by commas and round brackets. Commas act as conceptual separators and brackets assign prompt weights to each of the concepts (always kept at 1 here). There is also a one token long style descriptor that needs to be included at the beginning to invoke the trained style. Each concept descriptor contains a number, a quantity descriptor (few or many), the word associated with the element itself, and a colour.
@@
The training data was created to cover various possible floor plan designs from different building shapes up to buildings that have special arrangements like having no windows or a specific number of doors. The individual symbologies (like doors or living rooms) are also trained through images that only feature these symbols in several orientations.
@@
We fine-tuned a Stable Diffusion model v2.1 for each encoding $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ using Textual Inversions gal2022textual as preliminary trials showed that it offers fastest training time at equal result performance for our use case in comparison to hu2021lora and Ruiz_2023_CVPR .
3 Experiments
3.1 Validation Methodology
To compare the performance of the different un-tuned and fined-tuned diffusion models we defined a large set of 43 test prompts that evaluate different features representing practical request as well as challenging situations for the diffusion models. The prompts classify into the following categories:
@@ 3.1 Validation Methodology
Second, we check if the fine-tuned models do not overfit on the training samples, by prompting for human faces expecting that we do not get floor plans.
@@
To see if the model can understand quantities we query for “few” (less than 7) and “many” (more than 6) windows/doors and rooms/kittchen/bathrooms.
@@
For each test prompt and diffusion model in $\mathcal{B}$, $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, $\mathcal{SRE}$ we generate 10 sample images. In order to further improve results of all generations, we also added the following negative prompt to each iteration: ”gradients, blurry, fuzzy borders, reflections, lighting”. For each sample image we evaluated manually whether or not (1 or 0) the feature requested in the prompt was addressed in the generated image. This process resulted in 2.150 samples images that we evaluated.
3.2 Discussion
Table 1 shows the results of our evaluation. The baseline out-of-the-box diffusion model $\mathcal{B}$ without fine tuning performs worst. Only 6 % of all generated floor plans were fully valid, while most floor plans were faulty in some aspect like that rooms are missing walls, symbols are partially un-recognizable or room types are not clear (see Fig. 8). Ignoring these flaws, the model generates out-of-the-box something of a floor plan with some symbols for windows and doors and rooms recognizable. This indicates that Stable Diffusion knows the style of floor plans, but is not able to recreate fully.
Table 1: Quantitative results of the experimental validation of the four trained models: ##\## Number of Experiments; $*{}^{*}$ Ignoring full validity
@@ 3.2 Discussion
Based on this out-of-the-box experience, one may get the impression that Stable Diffusion is unable to generate useful floor plans. Using fine-tuning changes that drastically as visible in the valid plan category. The tuned models mostly manage to display a full building instead of partial cut-outs. There is also recurring symbology that avoids visual artifacts and errors that render a plan difficult to interpret or even entirely unreadable. The model with the best performance here is $\mathcal{SE}$. It is the most visually simple model, without room colouring or complex door and window symbols. Demonstrating that the Simplify strategy helps generating valid models.
@@
The overfitting test results show that the fine-tuned models are not overfitted. Even the base-line models do not always process the overfitting query correctly and all tuned models are in the same margin of error.
@@
Model $\mathcal{SE}$ with semantic element encoding was also the model that performing best to the request to remove specific elements. Most other models perform bad in this category. This shows that here our semantic Encoding strategy helps the diffusion model to learn what windows and doors are. That this is not strictly associated to the colour is evidenced with the results for the recolouring prompts, which usually failed.
@@
Similarly, the results of the count prompts show that the models simply are not yet capable of counting, even after fine-tuning models with many examples with correct counts. The results for quantitative prompts are significantly better. This may partially be because we only have two categories (few/many) resulting in a 50 % chance that the models get it right. This happens for example for the baseline model $\mathcal{B}$, which usually creates incomplete cut-off plans and, in consequential error, gets all few windows/doors/rooms queries correct. Yet, again the semantic encoding model $\mathcal{SE}$ performs above the random draw chance.
@@
The prompts testing room arrangements like L-, C- or O-shapes work well enough across all models at least in comparison to many other queries. This shows that even the baseline model has some basic understanding of arrangements in the higher diffusion model layers. This still is improved significantly with fine-tuning. Here the models $\mathcal{R}$ and $\mathcal{SE}$ show the highest performance increase. This is likely because the walls and rooms are very uniform in colour, increasing the contrast between inside and outside, which is harder to differentiate for the multicoloured rooms. Which supports our Contrasting strategy.
4 Results and Future Directions
4.1 Results
@@
The results from the experiments show that fine-tuning diffusion models significantly improves validity of the generated floor plans from 6 % of the un-tuned model to 57 %–90 % of the fine-tuned models. Tuning the models directly with semantic element encoding $\mathcal{SE}$ is resulting in the best performance across all test prompts. This is proving the correctness of our three strategies Simplify, Encode and Contrast as this model has the best trade-off between them. It is simple in containing only lines. By semantically encoding doors and windows with colours we keep them in the model without reducing simplicity and creating images with high contrast with only 4 colours. Adding room colours is reducing this contrast, which is why the $\mathcal{SR}$ and $\mathcal{SRE}$ models are performing worse.
@@
Our results also show that diffusion models still have major limitations in generating floor plans to satisfy queried requirements. The best performing queries for quantities do not exceed a 68 % success rate. Queries around counts and recolouring usually fail completely. Examples of common errors are shown in Fig. 8. They can be summarized as:
  • Results show only partial floor plans
  • Requested elements are not contained in the generated result
  • Symbols and colour palettes are generated in incorrect context
@@
  • Contextual layout instructions are ignored (e. g. kitchen left of living room)
  • Requested counts of elements and rooms are only met by chance
These issues lead us to some future research directions that we identified.
@@
Figure 8: Examples of generation errors: (a) overlapping door-like symbols; (b) incorrect symbols and colours; (c) building front-view instead of floor-plan
4.2 Future Directions
4.2.1 Future Directions for image-based Diffusion Models
@@
Our experiments revealed several limitations of diffusion models in generating floor plans leading to several future research directions. The success rates of prompts in our experiments are far from optimal. This is one of the reasons why diffusion models usually generate multiple samples at once and let the user to choose. They do than within seconds, significantly outperforming evolutionary generative design approaches. The analysis of Midjourney queries in ploennigs2022ai showed that only 13.1 % of all prompts are upscaled and then refined over 6 iterations—not minimizing the popularity of the platform. Here good workflows that allow to quickly select and refine queries for computational design are important to enable users to quickly filter out bad samples as shown in our idealized workflow in Fig. 2.
@@
Furthermore, the huge body of ongoing research in diffusion models will lead to improved models soon. An important topic in this fundamental area is their capability to correctly count elements in the generated images Yang_2023_CVPR . But, this still requires that the diffusion model has a good understanding of the semantics of the elements in a floor plan. If it does not know what symbol represents a door it will not be able to count it correctly.
@@
This lack of explicit semantic understanding is main limitation of diffusion models. Our experiments showed that it is possible to tune diffusion models to improve the results, but it is not currently feasible to remove all limitations and requires further research in semantic diffusion model encodings. Related to that is the question of how we can decode the semantic encodings of the generated images. Approaches within the area of diffusion models are restoration models, that for example allow to repair faces in generated images wang2021gfpgan .
@@
In relation to that we have the open research question of transforming generated binary images to Building Information Models that we initially wanted as results of the workflow in Fig. 2. The benefit of our semantic encoding is that our strategies of Simplify, Encode, and Contrast also create cleaner images for BIM restoration approaches like pizarro2022automatic .
4.2.2 Future Directions for BIM-based Diffusion Models
@@
All these future research topics above address the limitation of image-based diffusion models in computational design. What we actually want are BIM models. This leaves us with the final essential question: Why not develop a diffusion model for BIM models and simplify the whole process?
@@
From the introduction section 1.2.4 we can derive some requirements for such a BIM-based diffusion model:
@@
BIM models easily fulfil the last requirement. Given their nature, they are well structured and semantically rich models. Unfortunately, the first two requirements are not easily applicable to BIM models. We can of course raster a BIM model in different resolutions to generate binary images. But, our results show that we lose the semantic information and the well-structured nature in the process and that it is hard to recreate that.
Our proposal is to research in BIM-based Diffusion Models that keep that semantic information and utilizes the hierarchical nature of BIM models. In a BIM model we have floors, zones, and objects that form a hierarchy. Now the only requirement we have to solve is the diffusability. For that we need a Neural Network representation of the BIM model that we can easily alter. Here we can utilize the fact that at the core of BIM models like IFC lies a graph representation like STEP. Representing BIM models in Graphs was proposed before for generative approaches like gan2022bim ; li2023automated . But, we propose to research diffusion models instead of generative models for generation, as they can be trained with additional tags to generate specific designs.
The closest current neural network representation for this are Graph Neural Networks (GNN) recently developed for e. g. time series forecasting ba2022automated . We propose the research into multi-layer GNN model that represents on each layer the respective information of a BIM model, like which floors exists and which are close to each other, which zones exists within these floors and how they connect via walls. Diffusion approaches for graphs capable of weakening or removing those connections in the graph were recently developed chamberlain2021grand .
The benefit of the approach is that we can directly generate 3D models utilizing the BIM structure. A challenge here are the data requirements for training. As discussed in the beginning, training a diffusion model usually requires huge amounts of data. Here, it may be one option to use algorithmic design approaches to generate training data like we did for the fine-tuning. Alternatively, we as research community may need to start an open-source initiative to collect openBIM datasets at large scale to enable this next level of BIM-based Diffusion Models.

### v2

@@ 5.1 Generation of the Training Data
In order to fine-tune Stable Diffusion to the different semantic encodings that we described, we implemented an algorithmic design heuristic in Python, which procedurally generates a desired number of floor plans, encodes them in each of our four different proposed styles $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$, and then generates a customized description prompt for each variant. This ensures that all four models are on even ground during the training phase. Fig. 6 shows procedurally generated example images for each of the proposed encodings.
Figure 7: A low-resolution example for procedural generation in style $\mathcal{SRE}$.
The floor-plan generation algorithm is based on random room locations that are grown into rectangular regions until their edges collide with the edges and corners of other regions. A visual representation of the algorithm for one of the styles is shown in Fig. 7 and a description of the entire algorithm in simplified pseudocode is shown in Algorithm 1. Line 1–6 and images 1 and 2 show the setting of image bounds and the random selection of initial room coordinates. The process of growing the rooms is given in line 8–14 and images 3–7. The rooms grow in each direction (left $R_{\text{L}}$, right $R_{\text{L}}$ , top $R_{\text{L}}$ , bottom $R_{\text{L}}$) until they either intersect with another room or the image boundary. In lines 16–24 and images 8–11 we add walls, doors, and windows. Doors and windows are placed according to the simple rule that outside walls will often have a door flanked by two windows, while interior walls will only have one door. However, for every step and every element there is a chance for random alterations, such as the deletion of rooms or walls to create non-rectangular building arrangements. During this process we also render one variant of the plan for each style. Once the final plans are created, they are automatically checked for obvious errors and either fixed or discarded if necessary. This creates a collection of valid plans, rendered in the four different styles and labelled automatically. This final step is shown in lines 25–27. Afterwards, there is also a manual process of discarding any erroneous plans that the automated validation step did not catch, and of labelling additional aspects that would be challenging to automatically recognize, such as the general shape of the building (e.g. l-shape, o-shape, etc.). The code is publicly available on github [^1] https://github.com/AI4SC/bim-diffusion-models.
1:procedure Procedural_Generation($N_{G},N_{R},X_{\text{max}},Y_{\text{max}}$)
2:     for $g=0\ldots N_{G}$ do
3:         $B_{\text{x}}=\textsc{rand}\!\left(X_{\text{max}}\right)$; $B_{\text{y}}=\textsc{rand}\!\left(Y_{\text{max}}\right)$triangleright Compute random image boundary
4:         for $i=0\ldots N_{R}$ do $\triangleright$ For all rooms $\mathfrak{R}$ with $|\mathfrak{R}|=N_{R}$
5:              $R^{i}_{\text{x}}=\textsc{rand}\!\left(B_{\text{x}}\right)$; $R^{i}_{\text{y}}=\textsc{rand}\!\left(B_{\text{y}}\right)$triangleright Compute random room centers
6:              $R^{i}_{\text{RT}}=\textsc{rand}\!\left(\{\text{room},\text{kitchen},\text{bath}\}\right)$triangleright Assign random room type
7:         $\triangleright$ Repeat until no room can grow anymore
8:              no_update = true $\triangleright$ Init stop condition
9:              for each $R^{i}\in\mathfrak{R}$ do $\triangleright$ For all rooms
10:                  if $R^{i}_{\text{R}}+R^{j}_{\text{L}}<R^{i}_{\text{x}}-R^{j}_{\text{x}}\quad\forall R^{j}\in\mathfrak{R}|R^{j}_{\text{x}}<R^{i}_{\text{x}}$ then $\triangleright$ No intersection with rooms right of $i$
11:                       if $R_{\text{x}}+R^{i}_{\text{R}}\leq B_{\text{x}}$  then $\triangleright$ No intersection with $X$ boundary
12:                           $R^{i}_{\text{R}}=R^{i}_{\text{R}}+1$triangleright Increment room size right of center in $X$
13:                           no_update = false $\triangleright$ Reset stop condition
14:                  $\triangleright$ Similarly increment room size left $R^{i}_{\text{L}}$, top $R^{i}_{\text{T}}$, bottom $R^{i}_{\text{B}}$ considering axis neighbours and bounds
15:         until no_update
16:         for each $\text{S}\in\mathcal{\{\mathcal{R},\mathcal{SR},\mathcal{SE},\mathcal{SRE}\}}$ do $\triangleright$ Interate trough all style variants
17:              for each $R\in\mathfrak{R}$ do
18:                  $\mathfrak{W}_{R}=\textsc{Draw\_Walls}\!\left(R,S\right)$triangleright Draw walls in style $S$
19:                  for each $W\in\mathfrak{W}_{R}$ do
20:                       $\mathfrak{D}_{W}=\textsc{Draw\_Doors}\!\left(W,S\right)$triangleright Draw doors in style $S$
[…]
22:                           $\mathfrak{N}_{W}=\textsc{Draw\_Windows}\!\left(W,S\right)$triangleright Draw windows in style $S$
23:                  if $\text{S}\in\mathcal{\{\mathcal{SR},\mathcal{SRE}\}}$ then
24:                       $R_{\text{c}}=\textsc{Color\_Room}\!\left(S,R\right)$triangleright Semantic color room in style $S$
[…]
Algorithm 1 Algorithm for the procedural generation of floor plans in four styles with $N_{G}$ sample number, $N_{R}$ number of rooms, maximum image width $X_{\text{max}}$ and height $Y_{\text{max}}$
@@
The generated prompts describe the generated floor plan and follows the common guidelines for Stable Diffusion prompts, specifically for automatic1111: Individual concepts are separated by commas and round brackets. Commas act as conceptual separators and brackets assign prompt weights to each of the concepts (always kept at 1 here). There is also a style descriptor that is exactly one token long and needs to be included at the beginning of the prompt, in order to invoke the trained style from the textual inversion. Each concept descriptor contains a number, a quantity descriptor (few or many), the word associated with the element itself, and a colour, as shown in Fig. 8.
@@
5.2 Structure of the Training Data
The training data was created to cover various possible floor plan designs from different building shapes up to buildings that have special arrangements, like having no windows or a specific number of doors. The individual symbologies (like doors or living rooms) are also trained through manually created images that only feature these symbols in several orientations.
@@ 5.2 Structure of the Training Data
5.3 Fine-Tuning the Models
We fine-tuned a Stable Diffusion model v2.1 for each encoding $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ using Textual Inversions [34] as preliminary trials showed that it offers fastest training time at equal result performance for our use case in comparison to [32] and [31].
With each batch of generated training data, we fine-tuned a Stable Diffusion model v2.1 for each encoding $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ using Textual Inversions [34], as preliminary explained in Section 3.2. Preliminary trials showed that it offers fastest training time at equal result performance for our use case in comparison to [32] and [31]. We utilized the automatic1111 platform with the DreamBooth extension to run the fine-tuning. The source model was the v2-1_768-ema-pruned checkpoint by StabilityAI. We used the recommended training parameters for tuning a new style for this model, with 100 training epochs, a 768x768 maximum resolution and a one-token instance prompt. The latter creates a style descriptor that we then use to specifically invoke our fine-tuning during image generation.
We created a model variant for each style. Every model was able to attain reasonable similarity to the trained style for simple prompts, as shown in Fig. 9. We will evaluate the precise performance of each model for our use case through quantitative quality metrics in the next section.
(a) $\mathcal{R}$
(b) $\mathcal{SR}$
(c) $\mathcal{SE}$
(d) $\mathcal{SRE}$
Figure 9: Successful example generations for each of the four retrained models.
6 Evaluation
6.1 Methodology of Validation
To compare the performance of the different un-tuned and fined-tuned diffusion models we defined a large set of 43 test prompts that represent various practical request as well as specific challenging situations for the diffusion models. These prompts are based on the structure of the training prompts, and were created such that they each invoke different aspects of that training. The prompts are classified into the following categories, according to which challenges they pose to the model:
@@ 6.1 Methodology of Validation
Second, we check if the fine-tuned models do not overfit on the training samples. We consider a model overfitted if it responds to a prompt asking for a face with a floor plan, as “face” is semantically far from floor plan, we assume that the model always responds with floor plans even if not asked for and thus does not differentiate the tokens well enough.
@@
To see if the model can understand quantities we query for “few” (less than 7) and “many” (more than 6) windows/doors and rooms/kitchens/bathrooms.
@@
For each test prompt and diffusion model in $\mathcal{B}$, $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, $\mathcal{SRE}$ we generate 10 sample images. To further improve results of all generations, we also added the following negative prompt to each iteration: "gradients, blurry, fuzzy borders, reflections, lighting". For each sample image we evaluated manually whether or not (1 or 0) the feature requested in the prompt was addressed in the generated image. This process resulted in 2.150 sample images that we manually evaluated.
6.2 Results
Table 1 shows the results of our evaluation. The baseline out-of-the-box diffusion model $\mathcal{B}$ without fine tuning performs the worst. Only 6 % of all generated floor plans were fully valid, while most floor plans were faulty in some aspect. Examples for faults are rooms that are missing walls, symbols that are partially unrecognizable or room types that are not clear (see Fig. 10).
Table 1: Quantitative results of the experimental validation of the four trained models with ##\## Number of Experiments
@@ 6.2 Results
Based on these results, Stable Diffusion is unable to generate useful floor plans before additional fine-tuning. Once finetuned these capabilities drastically change, as visible in the valid plan category. The tuned models mostly manage to display a full building instead of partial cut-outs. There is also recurring symbology that avoids visual artifacts and errors that render a plan difficult to interpret or even entirely unreadable. The model with the best performance here is $\mathcal{SE}$. It is the most visually simple model, without room colouring or complex door and window symbols. Demonstrating that the Simplify strategy helps generating valid models.
@@
The overfitting test results show that the fine-tuned models are not overfitted. Even the baseline models do not always process the overfitting query correctly and all tuned models are in the same margin of error.
@@
Model $\mathcal{SE}$ was also the model that responds best to requests to removal of specific elements. Most other models perform badly in this category. This shows that our semantic Encoding strategy helps the diffusion model learn what windows and doors are. That learning this encoding is not strictly associated with just colour is evidenced by the results of the recolouring prompts, which usually failed.
@@
Similarly, the results of the count prompts show that the models simply are not yet capable of counting, even after fine-tuning models with many examples with correct counts. The results for quantify prompts are significantly better. This may partially be because we only have two categories (few/many), resulting in a high likelihood for false positives. The baseline model $\mathcal{B}$ for example usually creates incomplete cut-off plans with few elements, and thus gets all few windows/doors/rooms queries correct. Yet, the semantic encoding model $\mathcal{SE}$ seems to again perform above the random draw chance.
@@
The prompts testing room arrangements like L-, C- or O-shapes work well enough across all models, at least in comparison to many other queries. This shows that even the baseline model has some basic understanding of arrangements in the higher diffusion model layers. This still is improved significantly with fine-tuning. Here the models $\mathcal{R}$ and $\mathcal{SE}$ show the highest performance increase. This is likely because the walls and rooms are very uniform in colour, increasing the contrast between inside and outside, which is harder to differentiate for the multicoloured rooms. This result partially supports our Contrasting strategy, although it shows that the three strategies can sometimes run counter to each other and thus require careful balancing.
6.3 Discussion
@@
The results from the experiments show that fine-tuning diffusion models significantly improves validity of the generated floor plans from 6 % of the untuned model to 57 %–90 % of the fine-tuned models. Tuning the models directly with semantic element encoding $\mathcal{SE}$ is resulting in the best performance across all test prompts. This is proving the correctness of our three strategies Simplify, Encode and Contrast, as this model has the best trade-off between them. It is simple in that it contains only lines. By semantically encoding doors and windows with colours we keep them in the model without reducing simplicity and creating images with high contrast that feature only black, white, grey and two colours. Adding room colours reduces this contrast, which is why the $\mathcal{SR}$ and $\mathcal{SRE}$ models are performing worse.
@@
Our results also show that diffusion models still have major limitations in generating floor plans to satisfy queried requirements. The best performing queries for quantities do not exceed a 68 % success rate. Queries around counts and recolouring usually fail completely. Examples of common errors in the context of our evaluation are shown in Fig. 10. They can be summarized as:
  • Results show only partial floor plans.
  • Requested elements are not contained in the generated result.
  • Symbols and colour palettes are generated in incorrect context.
@@
  • Contextual layout instructions are ignored (e. g. kitchen left of living room).
  • Requested counts of elements and rooms are only met by chance.
There are of course other more general issues in the current generation of diffusion models. The more semantically detailed we want to get, the more difficult it would be to train models, like for example for the handling of different door and window sub-types and their correct symbols. Many models also have inherent quality limitations, for example regarding image size. Models can also reinforce bad design principles, as the aesthetics of image compositions that are encoded within them do not necessarily correspond to good architecture. These issues point to many important future research directions that we will identify in the following section.
@@
Figure 10: Examples of generation errors: a) overlapping door-like symbols; b) incorrect symbols and colours; c) building front-view instead of floor-plan encodings of the generated images. Approaches within the area of diffusion models are restoration models, that for example can repair faces in generated images [38]
7 Future Directions
7.1 Image-based Diffusion Models
@@
Our experiments revealed several limitations of diffusion models in generating floor plans. The success rates of prompts in our experiments are far from optimal. This is one of the reasons why diffusion models usually generate multiple samples at once and let the user choose. The generation itself happens within seconds, significantly outperforming evolutionary generative design approaches. The analysis of Midjourney queries in [1] showed that only 13.1 % of all prompts are upscaled and then refined over 6 iterations—which does not seem to be an issue for the popularity of the platform. Good workflows that allow users to quickly select and refine queries for computational design are important, to enable quick filtering of bad samples as shown in our idealized workflow in Fig. 2.
@@
Furthermore, the huge body of ongoing research in diffusion models will lead to improved models soon. An important topic in this fundamental area is their capability to correctly count elements in the generated images [39]. But, this still requires that the diffusion model has a good understanding of the semantics of the elements in a floor plan. If it does not know what symbol represents a door it will not be able to count it correctly.
@@
This lack of explicit semantic understanding is a main limitation of diffusion models. Our experiments showed that it is possible to tune diffusion models to improve the results, but it is not currently feasible to remove all limitations and requires further research in semantic diffusion model encodings. Related to this is the question of how we can turn visual semantic encodings like we used in this paper back into valid plan symbols. Possible approaches within the area of diffusion models are restoration models, that for example can repair faces in generated images [38]. The models have a similar lack of semantic understanding about additional context, like the surrounding buildings or site planning. While our results in specifying specific shapes were promising, we did not evaluate the external environment.
@@
There is also the open research question of transforming generated binary images to Building Information Models, which we identified as the ideal result of generative workflows in Fig. 2. The benefit of our semantic encoding is that our strategies of Simplify, Encode, and Contrast also create cleaner images for BIM restoration approaches like [37].
7.2 BIM-based Diffusion Models
@@
All the future research topics mentioned so far address the limitations of image-based diffusion models in computational design. However, the ideal domain-specific solution would be models that directly operate on BIM data. This leaves us with the final essential question: Why not develop a diffusion model for BIM models and thus strip out any information loss incurred by data conversions in and out of the process?
@@
From our introduction we can derive some requirements for such a BIM-based diffusion model:
@@
BIM models easily fulfil the last requirement. Given their nature, they are well structured and semantically rich models. Unfortunately, the first two requirements are not easily applicable to BIM models. We can of course raster a BIM model in different resolutions to generate binary images. But our results show that we lose the semantic information and the well-structured nature in the process, and that these aspects are difficult to recreate.
Our proposal is to research BIM-based Diffusion Models that instead keep semantic information and utilize the hierarchical nature of BIM models. In a BIM model we have floors, zones, and objects that form a hierarchy. Now the only requirement we must solve is the diffusability. For that, we need a Neural Network representation of the BIM model that we can easily alter. Here we can utilize the fact that at the core of BIM models like IFC lies a graph representation like STEP. Representing BIM models in Graphs was proposed before for generative approaches like [40, 41]. The closest current neural network representations for such structures are Graph Neural Networks (GNN) recently developed for e. g. time series forecasting [42]. We propose research into multi-layer GNN models that represent on each layer the respective information layer of a BIM model, like which floors exists and which are close to each other, which zones exists within these floors and how they are connected via walls. Diffusion approaches for graphs capable of weakening or removing those connections in the graph were recently developed [43]. Also, early approaches for using diffusion models for polygonal floor plans exist [44, 45], but they lack the semantic information of BIM.
The benefit of this approach is that, utilizing the BIM structure, we could directly generate 3D models instead of just 2D plans. A challenge here are the data requirements for training. As discussed in the beginning, training a diffusion model usually requires huge amounts of data. Here, it may be one option to use algorithmic design approaches to generate training data like we did for the fine-tuning in this paper. Alternatively, we as a research community may need to start an open-source initiative to collect open BIM datasets at a large scale to enable this next level of BIM-based Diffusion Models. How our workflow would look with this kind of model, is shown in Fig. 11.